OpenStudy (josephreak):

Need help! (Vectors) If the angle between two vectors, m = [3,1] & n = [2,x] is pi/4. Find X I am using the formula (which I'm sure is the correct one) cos(theta) = (m crossproduct n)/|m| |n| I plug in but I am doing something wrong when it comes to simplifying, something, somewhere, goes wrong and it messes it all up. Thanks.

OpenStudy (phi):

I would use a dot product

OpenStudy (phi):

\[ m \cdot n = 2\cdot 3 + 1\cdot 1 = x+6 \] the magnitude of m is sqr(10) mag of n is \(\sqrt{x^2+4} \) using \[ m \cdot n = |m| |n| \cos \frac{\pi}{4}\\ \frac{m \cdot n}{|m| |n|}= \frac{1}{\sqrt{2} } \]

OpenStudy (josephreak):

There's a typo in my post, I meant dot product* Also, I solved it.. it turned out I was doing everything completely right... but I stupidly forgot to properly expand the (x+6)^2 lol thanks for your help though!